# Mihai Dascalub, Traian Rebedeab,f

### v2

@@ 7 Conclusions
We present a general training and evaluation recipe which we expect to work on different, more powerful LLMs as well (e.g., models larger than 7-8B, Mixtral121212www.mistral.ai/news/mixtral-of-experts/ models) as well as for other languages. We firmly believe that this is just the first step towards building and adapting LLMs for Romanian for both research and industry applications.
Pretraining on additional cleaner data, alignment with human preference Ziegler et al. (2019); Rafailov et al. (2024); Dong et al. (2023), validating existing translations, and adding more instructions and conversations are just a few improvements that should increase the quality of RoLLMs.
@@ Acknowledgements
The universities contributed with researchers who worked pro-bono (POLITEHNICA Bucharest and the University of Bucharest) and computing power for training the models (POLITEHNICA Bucharest). Synchronization and collaboration took place within a project131313www.ilds.ro/llm-for-romanian carried out in the Institute of Logic and Data Science with funding from BRD Groupe Societe Generale.
We thank Alin Stefanescu and Laurentiu Leustean for participating in various discussions regarding this project.
Parts of this work were presented in  Masala et al. (2024). We improve upon previous work by training and evaluating more models, adding downstream tasks evaluation, the novel RoCulturaBench and more general training and evaluation tools [^14] https://github.com/OpenLLM-Ro.
Limitations
Our approach for adapting existing LLMs to a new language (i.e., Romanian) has some notable limitations. One significant drawback is the reliance on translated (mainly single-turn) instructions for the instruction fine-tuning step. Subtle nuances might not be captured by the translation process, a process that might introduce inaccuracies that decrease the model’s performance. Additionally, our models lack any kind of safeguard or moderation mechanism and are not specifically aligned with human values. That means that our models could generate inappropriate, racist, toxic, dangerous, or even illegal content, raising ethical and practical concerns.
@@ A.1 MT-Bench
Llama-2-7b-chat (ro_prompted) 1.90 1.65 1.32 0.95 0.30 0.75 1.75 1.05
RoLlama2-7b-Instruct 4.40 4.45 3.17 1.65 2.25 4.00 4.25 5.20
Mistral-7B-Instruct-v0.2 (ro_prompted) 4.45 5.00 3.42 3.77 4.35 4.90 5.25 7.30
RoMistral-Instruct 5.50 4.45 4.32 3.32 3.30 5.70 5.90 6.65
Llama-3-8B-Instruct (ro_prompted) 7.00 6.45 4.85 3.60 5.25 4.47 5.90 6.85
RoLlama3-8b-Instruct 5.85 5.90 3.22 4.00 3.45 6.70 5.85 6.85
gemma-1.1-7b-it (ro_prompted) 5.10 4.80 3.85 3.05 4.05 5.60 4.95 5.60
RoGemma-7b-Instruct 5.40 5.50 2.68 4.50 3.15 6.16 5.75 5.25
Table 8: Performance of instruct models on each category of Romanian MT-Bench. W - Writing, Rp - Roleplay, R - Reasoning, M - Math, Code - Coding, Ext - Extraction, H - Humanities. Bold denotes the best result for each category.
@@ A.2 RoCulturaBench
Llama-2-7b-chat (ro_prompted) 1.72 2.25 0.53
RoLlama2-7b-Instruct 3.43 4.78 1.35
Mistral-7B-Instruct-v0.2 (ro_prompted) 3.75 4.95 1.20
RoMistral-Instruct 3.17 4.40 1.23
Llama-3-8B-Instruct (ro_prompted) 4.61 6.14 1.53
RoLlama3-8b-Instruct 3.83 5.06 1.23
gemma-1.1-7b-it (ro_prompted) 3.22 4.35 1.13
RoGemma-7b-Instruct 3.47 4.56 1.09
Table 9: Performance of Instruct models on RoCulturaBench. Bold denotes the best result for each category.
@@
After manually inspecting part of the results, we found that RoLLMs tend to have a more rigid answer. The other models have a more conversational tone, and they add a lot of details (that are correct but not specifically requested by the prompt), which is more in line with the reference answers and general human preference.
On average, answers generated by these models are over 60% longer than answers generated by RoLLMs and only 10% shorter than the reference answers. This sometimes rigid tone of RoLLMs stems mostly from the instruction dataset that was used for fine-tuning and the lack of human preference alignment mechanisms. A more diverse set of instruction coupled with human-alignment methods will alleviate a lot of this tone, making the models more chatty.
@@
Model W R A G H C&S Sport C&T S C
Llama-2-7b-chat (ro_prompted) 0.0 1.4 3.0 2.8 1.5 2.3 1.6 1.2 1.4 2.2
RoLlama2-7b-Instruct 3.1 4.0 5.7 5.2 3.3 2.6 3.4 2.4 2.6 2.6
Mistral-7B-Instruct-v0.2 (ro_prompted) 3.9 3.8 6.2 4.2 3.4 3.2 3.1 3.0 4.4 2.8
RoMistral-Instruct 3.1 4.1 4.8 3.3 3.8 2.5 2.4 2.3 3.6 2.2
Llama-3-8B-Instruct (ro_prompted) 5.4 5.6 6.6 4.9 4.3 3.8 4.0 4.1 5.2 3.3
RoLlama3-8b-Instruct 4.7 5.3 5.7 4.5 3.2 3.0 3.1 3.4 3.1 2.0
gemma-1.1-7b-it (ro_prompted) 3.2 4.4 5.5 3.7 3.2 2.1 2.5 2.5 3.1 2.0
RoGemma-7b-Instruct 3.3 4.6 5.9 4.5 2.7 2.8 2.3 3.4 3.8 1.9
Gpt-3.5 5.5 6.6 7.1 6.7 5.9 4.5 4.4 6.5 5.2 4.6
Table 10: Performance of Instruct models on each category of RoCulturaBench. W - Writing, R - Roles, A - Argumentation, G - Geography, H - History, C&S - Culture & Science, C&T - Customs & Traditions, S - Stereotypes, C - Celebrity. Bold denotes the best result for each category.
@@ Appendix E Computational Resources
Training and evaluation were performed on up to six NVIDIA A100s with 80GB of memory. Continual pre-training on 5% of the data using a sequence length of 512 for one epoch used 320 GPU hours, 468 GPU hours were consumed for a sequence length of 1024, while for a sequence length of 2048, 708 GPU hours were needed. Instruction fine-tuning of RoLlama2 with a sequence length of 1024 for one epoch required 52 GPU hours.
The evaluation was performed on a single A100 GPU at a time.

### v4

@@ 7 Conclusions
We present a general training and evaluation recipe which we expect to work on different, more powerful LLMs as well (e.g., models larger than 7-8B, Mixtral141414www.mistral.ai/news/mixtral-of-experts/ models) as well as for other languages. We firmly believe that this is just the first step towards building and adapting LLMs for Romanian for both research and industry applications.
Pretraining on additional cleaner data, validating existing translations, and adding more instructions, conversations and human preference data are just a few improvements to increase the quality of RoLLMs.
Limitations
Our approach for adapting existing LLMs to a new language (i.e., Romanian) has some notable limitations. One significant drawback is the reliance on translated (mainly single-turn) instructions for the instruction fine-tuning step. Subtle nuances might not be captured by the translation process, a process that might introduce inaccuracies that decrease the model’s performance. Additionally, our models lack any kind of safeguard or moderation mechanism, meaning that they could generate inappropriate, racist, toxic, dangerous, or even illegal content, raising ethical and practical concerns.
@@ Acknowledgements
Synchronization and collaboration for this work took place within a project151515www.ilds.ro/llm-for-romanian carried out in the Institute of Logic and Data Science with funding from BRD Groupe Societe Generale. Part of this work was funded by the EU’s NextGenerationEU instrument through the National Recovery and Resilience Plan of Romania – Pillar III-C9-I8, managed by the Ministry of Research, Innovation and Digitalization, within the project entitled Measuring Tragedy: Geographical Diffusion, Comparative Morphology, and Computational Analysis of European Tragic Form (METRA), contract no. 760249/28.12.2023, code CF 163/31.07.2023. We thank Alin Stefanescu and Laurentiu Leustean for participating in various discussions regarding this project.
@@ A.1 MT-Bench
Llama-2-7b-chat (ro_prompted) 1.95 1.15 0.70 1.00 0.25 1.35 1.35 0.90
RoLlama2-7b-Instruct-DPO 5.60 6.25 4.35 2.30 2.25 3.95 5.35 6.80
Mistral-7B-Instruct-v0.2 (ro_prompted) 4.70 5.00 5.40 3.80 4.85 3.65 5.50 7.30
RoMistral-7b-Instruct-DPO 6.75 6.50 6.90 4.50 3.35 5.60 5.85 7.60
Llama-3-8B-Instruct (ro_prompted) 7.20 6.25 7.55 4.30 5.40 4.70 5.70 6.60
RoLlama3-8b-Instruct-DPO 6.90 6.60 7.25 4.80 3.35 4.55 6.60 6.90
Llama-3.1-8B-Instruct (ro_prompted) 6.55 4.95 7.35 5.50 5.75 3.55 5.25 6.63
RoLlama3.1-8b-Instruct-DPO 6.85 6.65 7.90 6.25 3.95 4.55 6.00 7.55
gemma-1.1-7b-it (ro_prompted) 5.45 5.00 5.75 3.45 3.80 4.05 5.00 6.15
RoGemma-7b-Instruct-DPO 5.90 5.95 6.35 5.35 3.55 4.55 5.65 6.60
Table 9: Performance on each category of Romanian MT-Bench. W - Writing, Rp - Roleplay, R - Reasoning, M - Math, Code - Coding, Ext - Extraction, H - Humanities. Bold denotes the best result for each category.
@@ A.2 RoCulturaBench
Llama-2-7b-chat (ro_prompted) 1.21 2.42 1.21
RoLlama2-7b-Instruct-DPO 4.80 5.95 1.15
Mistral-7B-Instruct-v0.2 (ro_prompted) 3.68 4.48 0.80
RoMistral-7b-Instruct-DPO 4.72 5.88 1.16
Llama-3-8B-Instruct (ro_prompted) 4.62 5.60 0.98
RoLlama3-8b-Instruct-DPO 4.40 5.38 0.98
Llama-3.1-8B-Instruct (ro_prompted) 3.54 4.50 0.96
RoLlama3.1-8b-Instruct-DPO 4.42 5.51 1.09
gemma-1.1-7b-it (ro_prompted) 3.38 4.07 0.69
RoGemma-7b-Instruct-DPO 3.94 4.71 0.77
Table 10: RoCulturaBench performance with and without reference answers.
@@
Model W R A G H C&S Sp C&T S C
Llama-2-7b-chat (ro_prompted) 0.0 1.5 1.5 1.6 2.2 1.7 0.3 0.6 1.4 1.3
RoLlama2-7b-Instruct-DPO 5.4 6.0 6.0 5.2 4.8 3.9 3.4 4.7 4.3 4.3
Mistral-7B-Instruct-v0.2 (ro_prompted) 4.1 3.3 5.6 3.6 3.7 3.5 3.1 2.7 4.5 2.7
RoMistral-7b-Instruct-DPO 5.6 5.7 7.1 5.5 3.6 2.9 4.3 4.1 4.9 3.5
Llama-3-8B-Instruct (ro_prompted) 5.4 5.8 6.7 4.9 3.8 3.5 3.3 4.2 5.0 3.6
RoLlama3-8b-Instruct-DPO 5.3 5.6 5.8 5.5 3.7 2.9 3.8 4.0 4.6 2.8
Llama-3.1-8B-Instruct (ro_prompted) 4.1 4.3 5.3 3.6 2.7 2.6 3.0 3.1 4.0 2.7
RoLlama3.1-8b-Instruct-DPO 4.9 5.7 6.8 5.2 3.5 3.4 3.1 3.7 4.7 3.2
gemma-1.1-7b-it (ro_prompted) 3.9 4.8 5.6 3.6 3.2 2.3 2.5 2.7 3.4 1.8
RoGemma-7b-Instruct-DPO 5.0 5.0 6.5 3.8 3.4 2.2 3.0 3.5 4.9 2.1
Gpt-3.5 6.0 7.0 7.8 7.1 5.7 4.8 4.9 6.6 5.8 5.2
Table 11: Performance on each category of RoCulturaBench. W - Writing, R - Roles, A - Argumentation, G - Geography, H - History, C&S - Culture & Science, C&T - Customs & Traditions, Sp - Sports, S - Stereotypes, C - Celebrity. Bold denotes the best result for each category.
@@ Appendix E Computational Resources
Pre-training was performed on up to six NVIDIA A100s with 80GB of memory. Continual pre-training on 5% of the data using a sequence length of 512 for one epoch used 320 GPU hours, 468 GPU hours were consumed for a sequence length of 1024, while for a sequence length of 2048, 708 GPU hours were needed. Instruction fine-tuning of RoLlama2 with a sequence length of 1024 for one epoch required 120 GPU hours, while performing DPO required 30 GPU minutes for one epoch. Evaluations were performed on a single A100 GPU at a time.